OpenStudy (anonymous):

Rationalize the denominator: cubic root of 2y^4/6x^4

OpenStudy (anonymous):

\[\sqrt[3]{2y ^{4}/6x ^{4}}\]

OpenStudy (anonymous):

My instinct is to go down instead of making it \[x ^{64}\] because that would look dumb, unless it's correct but I've never had to decrease... do I just multiply by a negative exponent?

OpenStudy (anonymous):

=cube root (2y^4)/cube root(6x^4) =cube root(2y^4)*cuberoot(6x^4)/cube root(6x^4)*cube root(6x^4) =cube root(2y^4)*cuberoot(6x^4)/6x^4 basically separate top and bottom roots multiply by conjugates fix ugly maths

jhonyy9 (jhonyy9):

cuberoot(y/3x)^4 =(y/3x)cuberoot(y/3x)=(ycuberoot3xy)/3x

jhonyy9 (jhonyy9):

2/6 simplifie by 2 and get 1/3

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

Hmm, well that is handy, didn't know oyu could simplify it but what would the correct conjugate be?

OpenStudy (anonymous):

you*

jhonyy9 (jhonyy9):

good luck bye

jhonyy9 (jhonyy9):

1/sqrt3 =(sqrt3)/3 ---like example

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

No actually, I'm confused... can you just tell me what I need to multiply by to get rid of the radical in the denominator?

jhonyy9 (jhonyy9):

1/sqrt3 so when there is radical in denominator multiplie numerator and denominator by conjugate so in this case by sqrt3 and in this way you get in denominator 3 and in numerator sqrt3

jhonyy9 (jhonyy9):

because sqrt3 * sqrt3 =3

OpenStudy (anonymous):

I know... but what would the conjugate be. I know how to do all this already i get it I just need to know if the conjugate I am using is correct...

jhonyy9 (jhonyy9):

your answer not is right

jhonyy9 (jhonyy9):

check my answer step by step

razor99 (razor99):

we have to simplify that.

OpenStudy (anonymous):

Ok look, the denominator is cubic root of 6x^4 .... so what do I multiply it by to take a perfect cubic root?

OpenStudy (anonymous):

sorry that is wrong

OpenStudy (anonymous):

\[\sqrt[3]{36x^2}\]

OpenStudy (saifoo.khan):

If sat is here, then there was no need of calling me! :)

OpenStudy (anonymous):

then you will get in the denominator a perfect cube, namely \[\sqrt[3]{216x^6}\]

OpenStudy (anonymous):

He just got here thanks anyway

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

you have \[\sqrt[3]{6x^4}\] so to make it a perfect cube, you need the exponent on the x to be 6, and you have to multiply by 6^2=36 to make the constnant a cube

OpenStudy (anonymous):

Wait why does the exponent for x have to be 6?

OpenStudy (anonymous):

because you have a cube root so you need a number divisible by 3

OpenStudy (anonymous):

i mean an exponent divisible by 3

OpenStudy (anonymous):

\[\sqrt[3]{216x^6}=6x^2\]

OpenStudy (anonymous):

So that would come out as \[6x ^{2}\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ah thank you that is ALL I needed, I guess nobody understood what I meant. Thank you very much

OpenStudy (anonymous):

yw